OpenStudy (anonymous):

Please help I'm lost and not sure if my answer is correct X-1= 6 ---- --- X. 8 Then in the end I got 13x = -8 13. 13 So X=0.61

pooja195 (pooja195):

Hi Welcome to openstudy :D x-1=6 is the original problem correct?

OpenStudy (anonymous):

The original problem is X-1= 6 X. 8

OpenStudy (mathmale):

You might be better off drawing the problem in the Draw utility:

OpenStudy (anonymous):

Yes correct

pooja195 (pooja195):

\[\huge~\rm~\bf~\color{purple}{\frac{ x-1 }{ x }=\frac{ 6 }{ 8}}\] Use cross multiplcation :-) Do you know how to do cross multiplcation?

OpenStudy (anonymous):

Yeah I got 6x=8x-8

OpenStudy (anonymous):

Is that right?

pooja195 (pooja195):

Yep go on and isolate x :-)

OpenStudy (anonymous):

Ok then I did +8x +8x So then I got 13x=-8

OpenStudy (anonymous):

13x. -8 --- =. --- 13. 13

pooja195 (pooja195):

Hmm not quite... \[\huge~\rm~\bf~\color{red}{6x-8x=8x-8x-8}\] We would subtract 8x from both sides because we want to ISOLATE x in other words get it by itself so... 6-8=?

OpenStudy (anonymous):

oh wow your right I didn't notice that

pooja195 (pooja195):

:3 now you can try to isolate x :)

OpenStudy (anonymous):

So then my final answer was X= -4 since there is a -8

OpenStudy (anonymous):

Correct? :)

pooja195 (pooja195):

Hmm not quite... \[\huge~\rm~\bf~\color{orange}{6x-8x=8x-8x-8}\] \[\huge~\rm~\bf~\color{orange}{-2x=-8}\] \[\huge~\rm~\bf~\color{red}{negative/negative=+ }\]

pooja195 (pooja195):

Does that make sense?

OpenStudy (anonymous):

Oh my lol I didn't notice that either haha So then I get -2x. -8 ---- = ---- -2. -2 So X= +4

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

Pooja you there? :(

pooja195 (pooja195):

Oops sorry >.< this lag >_> yes that's correct ^_^ x=4

OpenStudy (anonymous):

Yay !!! Thank you so much !!! :)